OpenStudy (anonymous):

at her store, stella sells clothing 60% more than she paid. during a weekend sale, stella sold clothing for 12.5% off what she would normally charge. her revenues, including the 6% florida sales tax, totaled $29,680. how much did stella pay, in dollars, for clothing sold that weekend?

OpenStudy (anonymous):

if we assume A as the price of weekend for each cloth she should pay 6% of A for tax. so A-0.06*A=29.68$ and then A=31.57$ .it means she earns 31.57$ per cloth at weekend and she should pay 6% of it for tax which is (0.06*31.57)=1.89$

OpenStudy (anonymous):

how did you get 31.57 for a

OpenStudy (anonymous):

if we assume A as the price of weekend for each cloth she should pay 6% of A for tax. so A-0.06*A=29.68$ then 0.94*A=2968

OpenStudy (anonymous):

but you know that isnt rite becuase it is 29.68 its 29,680

OpenStudy (anonymous):

ok, A=31,574$ right?

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

thats wron

OpenStudy (anonymous):

because when you plug it in it does not work out

OpenStudy (anonymous):

whatever ignore me

OpenStudy (anonymous):

!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!